At time t = 0, the displacement x(0) of the block in a linear oscillator like that of Fig. 3

is -8.50 cm. The block’s velocity v(0) then is -0.93 m/s, and its acceleration a(0) is 45.0

m/s2 . Determine,

(i) the angular frequency of this system,

(ii) phase constant and

(iii) amplitude of this system. [3,5,3]


1
Expert's answer
2021-05-21T10:42:35-0400

Solution.

"t=0;"

"x(0)=-8.50cm=-0.085m;"

"v(0)=-0.93m\/s;"

"a(0)=45.0m\/s^2;"

"i) x(t)=Acos(\omega t+\phi_0);"

"v(t)=-A\omega sin(\omega t+\phi_0);"

"a(t)=-A\omega ^2cos(\omega t+\phi_0);"

"-0.085=Acos\phi_0\implies cos\phi_0=-\dfrac{0.085}{A};"

"-0.93=-A\omega sin\phi_0;"

"45=-A\omega^2cos\phi_0=\omega^2\sdot 0.085\implies \omega^2=529.4; \omega=23rad/s;"

"ii) -0.93=-23 Asin\phi_0;"

"-0.085=Acos\phi_0";

"10.94=-23tan\phi_0\implies tan\phi_0=-0.475;\phi_0=2.7rad;"

"iii) A=\dfrac{-0.085}{cos\phi_0}=0.094m;"

Answer:"i) \omega=23rad/s;"

"ii)\phi_0=2.7rad;"

"iii)A=0.094m" .